OpenStudy (anonymous):

Solve the equation 3x^2 + 24x = 0 for x.Select one of the options below as your answer:A. -8, 0B. 8, 0C. -8D. 8

OpenStudy (imtant):

Factorize it to get x values. So it is 3x(x+8)=0 3x=0, x=0/3, x=0 for (x+8)=0, x=-8 the answer is A

OpenStudy (anonymous):

thank you

OpenStudy (imtant):

Wc.:)